# Investigations of MWISP Clumps: 13CO Clump Source Catalogs and Physical Properties

### v1

@@ 5.3.1 Cross-Survey Comparison of Clump Scaling Relations
Figure 11 presents a detailed comparison of molecular clumps scaling relations across catalogs, which can be directly assessed against Larson’s classical scaling laws (Larson, 1981) that have served as key references in molecular cloud dynamics.
The $M$ - $R_{\text{eq}}$ relation for molecular clumps (Figure 11a) exhibits a power-law index of $\alpha_{0.9}=2.229\pm 0.002$ ($P=0.97$), with CHIMPS ($\alpha_{0.9}=2.383$) and SEDIGISM ($\alpha_{0.9}=2.28$) displaying nearly identical slopes. These slopes, significantly steeper than the classical $M\propto R^{2.0}$ relation, indicate that surface density increases with size rather than remaining constant (Solomon et al., 1987; Lombardi et al., 2010). This departure from constant surface density spans multiple scales may reflect hierarchical structure formation where material becomes increasingly concentrated at larger scales (Heitsch et al., 2009; Ballesteros-Paredes et al., 2011; Vázquez-Semadeni et al., 2017), potentially arising from the intrinsic structure of supersonic, intermittent turbulent flows (Roman-Duval et al., 2010).
@@
Figure 11(f) presents the $\sigma_{v}$ - $R_{\mathrm{eq}}\Sigma$ relationship, yielding a power-law index of $\alpha_{0.9}=0.303\pm 0.001$ with strong correlation ($P=0.83$) for MWISP clumps, paralleling earlier MWISP datasets (Feng et al., 2024). Both CHIMPS ($\alpha_{0.9}=0.383$, $P=0.69$) and SEDIGISM ($\alpha_{0.9}=0.412$, $P=0.76$) display slightly higher indices compared to MWISP. This relation demonstrates superior statistical robustness compared to the $\sigma_{v}$ - $R_{\mathrm{eq}}$ relation ($P=0.83$ versus $P=0.73$) (Heyer & Dame, 2015; Miville-Deschênes et al., 2017). For clumps in virial equilibrium, the virial theorem predicts $\sigma_{v}\propto(R\times\Sigma)^{0.5}$. The observed exponents remain below the virial prediction, indicating sub-virial states where turbulent support scales more weakly with gravitational potential than expected in idealized equilibrium models.
While Larson’s original relations were based on relatively limited samples and the technologies available over 40 years ago (Larson, 1981), the comprehensive coverage and high precision of modern surveys like MWISP reveal nuances in molecular clump structure and dynamics that could not be captured previously (e.g. Solomon et al., 1987; Heyer et al., 2009; Lombardi et al., 2010; Ballesteros-Paredes et al., 2018). The MWISP survey’s unique characteristics, including large-scale coverage, an extensive sample, and uniform detection and distance analysis methods, make it well-suited for analyzing correlations between parameters. This minimized systematic effects across diverse galactic environments allow observed scaling relations to reflect more genuine physical processes. Rather than challenging the importance of Larson’s pioneering work, these results represent a natural advancement in our understanding of interstellar matter distribution enabled by modern observational capabilities.
Figure 12: Scaling relations between physical parameters of MWISP. Panel shows: (a) $M$-$R_{\rm eq}$; lightcoral dotted lines indicate constant $\Sigma$ = 1, 10, 40, 100, 1000 Mpc-2 and lime dotted lines indicate constant $n({\rm H_{2}})$ = 10, 100, 1000, 10000 cm-3; (b) $n({\rm H_{2}})$-$R_{\rm eq}$; (c) $\alpha_{\rm vir}$-$R_{\rm eq}$; (d) $\sigma_{v}$-$R_{\rm eq}$; (e) $\sigma_{v}$-$M$; (f) $\sigma_{v}$-$R_{\rm eq}\Sigma$; (g) $N(^{13}\mathrm{CO})$-$R_{\rm eq}$; (h) $\alpha_{\rm vir}$-$N(^{13}\mathrm{CO})$; and (i) $\sigma_{v}$-$N(^{13}\mathrm{CO})$. The distributions are shown for two subsets: MWISP complete clumps (gray points with cyan contours) and MWISP clumps with maser-determined distances (CMD). Among the CMD subset, CMD-C, indicated by colored points with blue contours, represents clumps in the close vicinity of masers, while CMD-H, marked by red contours, represents clumps that host masers. For each relation, the best-fit power-law index and corresponding $P$ are indicated in the legend.
@@ 5.3.2 Impact of Star Formation Activity on Clump Scaling Relations
Figure 12 contrasts scaling relations between the complete MWISP sample and a subset of clumps with maser-determined distances (CMD), further subdivided into clumps close to maser emissions (CMD-C, 1,566 sources) and clumps directly hosting masers (CMD-H, 152 sources). These subsets represent different levels of star formation activity: CMD-H captures clumps with ongoing maser excitation processes indicative of active star formation (e.g. Menten, 1991; Cragg et al., 2005; Breen et al., 2010; Urquhart et al., 2011b), CMD-C represents regions influenced by nearby star formation activity, while the complete sample encompasses diverse star formation environments ranging from quiescent to active phases.
@@
The $\alpha_{\rm vir}$-$R_{\rm eq}$ relation (Figure 12c) displays negative correlations, with changes in slope from the complete sample ($\alpha_{0.9}=-0.728$) to CMD-C ($\alpha_{0.9}=-0.69$, P=0.56P=-0.56) to CMD-H ($\alpha_{0.9}=-0.582$, P=0.59P=-0.59). This progressive flattening of the power-law slope, combined with the rightward and downward displacement of CMD-H distribution, indicates that maser-hosting clumps are systematically larger and more gravitationally bound. The flattening trend suggests that virial parameters become less sensitive to radius variations in active star formation environments, possibly reflecting stellar feedback processes that provide scale-independent support mechanisms across different clump sizes, partially balancing gravitational binding with non-thermal support to maintain relatively uniform dynamical states in these regions.
Panels (d), (e), and (f) in Figure 12 demonstrate changes in velocity dispersion scaling relations across different star formation environments. In the $\sigma_{v}$$R_{\rm eq}$ relation (panel d), both maser-associated populations display flatter slopes ($\alpha_{0.9}=0.152,0.161$, $P=0.27,0.41$) compared to the complete sample ($\alpha_{0.9}=0.396$). This pattern continues in the $\sigma_{v}$-$M$ relation (panel e), where power-law indices decrease from the complete sample ($\alpha_{0.9}=0.172$) to CMD-C ($\alpha_{0.9}=0.151$, $P=0.63$) and CMD-H ($\alpha_{0.9}=0.124$, $P=0.55$). Similarly, the $\sigma_{v}$-$R_{\rm eq}\Sigma$ relation (panel f) shows a decline from the complete sample ($\alpha_{0.9}=0.303$) to CMD-C ($\alpha_{0.9}=0.287$, $P=0.72$) and CMD-H ($\alpha_{0.9}=0.261$, $P=0.74$). The overall flattening of these scaling relations indicates that velocity dispersion becomes less sensitive to physical parameters in active star formation environments. The relatively small differences between CMD-C and CMD-H coefficients suggest that these modifications to turbulent properties occur in association with star formation activity, irrespective of whether clumps directly host masers.
@@
The $\sigma_{v}$-$N(^{13}\mathrm{CO})$ relation (Figure 12i) shows substantial reduction in slopes from the complete sample ($\alpha_{0.9}=0.801$, $P=0.82$) to identical values for both CMD-C ($\alpha_{0.9}=0.434$, $P=0.63$) and CMD-H ($\alpha_{0.9}=0.429$, $P=0.72$). This relation provides particularly valuable insights as both parameters are distance-measurement-independent quantities, eliminating potential biases from distance uncertainties. The power-law indices demonstrate higher density regions in general clumps maintain proportionally higher turbulent motions. The identical slops between CMD-C and CMD-H populations indicate that these relationships converge to similar characteristics in maser-associated environments, suggesting that turbulent properties may be regulated by similar physical processes (Heyer & Brunt, 2004) in maser-associated regions.
@@ 6 Summary
We present an equivalent global detection scheme that significantly reduces computational resource requirements while preserving detection integrity for processing large astronomical datasets. By integrating this scheme with FacetClumps, we successfully extract 71,661 molecular clumps across the MWISP Phase I survey’s high-resolution 13CO data spanning 2310 deg2. The designed automatic hierarchical distance decision method, using signal regions as fundamental objects, effectively resolves the kinematic distance ambiguity problem, offering reliable distance measurements for 97.94% of clumps.
The MWISP clumps have median values of 225.78 M in mass, 1.72 pc in radius, 0.4 km s-1 in velocity dispersion, 25.53 M pc-2 in surface density, and 214.69 cm-3 in volume density. With a median virial parameter of $\alpha_{\text{vir}}=1.44$, these structures are predominantly gravitationally bound, as 65.3% satisfy the criterion $\alpha_{\text{vir}}<2$ and account for 96.3% of the total statistical mass. The MWISP clump catalog uniquely traces intermediate-mass and intermediate-size structures with narrow velocity dispersions and low densities. These clumps, characterized by extended, diffuse molecular distributions, represent an important component of the Galactic molecular clump population.
@@
    The 13CO column density - radius relation ($\bar{N}(^{13}\mathrm{CO})\propto R^{0.296}$) shows gradual column density increase with clump size.
@@
The MWISP clump catalogs feature broad coverage, effectively capturing the critical transition from diffuse molecular gas to gravitationally collapsing structures. Despite inherent intrinsic survey limitations, completeness limits, and distance effects, these catalogs achieve exceptional statistical robustness and represent the most extensive study of molecular clumps to date. This unique characteristic establishes these catalogs the basis for a series of investigations, including: (1) analyzing the distributions of clump parameters across the Galactic disk and their trends with Galactocentric distance (Sun et al., 2024b); (2) identifying filaments using clump masks, directional, and positional properties (Jiang et al., 2025); (3) integrating multi-wavelength observations to accurately classify the evolutionary stages of star-forming regions (Urquhart et al., 2022) and derive their associated scaling relations; and numerous additional applications. These investigations will significantly advance our understanding of the complex interplay between molecular clumps, star formation, and Galactic evolution.
@@ Acknowledgements
We are grateful to the anonymous referee for the invaluable insights and comments, which enabled us to refine and enhance this work. This work is supported by the National Key R&D Program of China (grant No. 2023YFA1608001) and the National Natural Science Foundation of China (grants No. U2031202). This research make use of the data from the Milky Way Imaging Scroll Painting (MWISP) project, which is a multiline survey in 12CO/13CO/C18O along the northern galactic plane with PMO-13.7m telescope. We are grateful to all the members of the MWISP working group, particularly the staff members at PMO 13.7m telescope, for their long-term support. MWISP was sponsored by the National Key R&D Program of China with grants 2023YFA1608000, 2017YFA0402701, and the CAS Key Research Program of Frontier Sciences with grant QYZDJSSW-SLH047.
@@ Appendix A Configuration Parameters
Table 2: Parameters of FacetClumps along with their explanations and values.
@@
FacetClumps.RMS The noise rms or S/N of the data. 1 (S/N data)
@@
FacetClumps.SWindow The scale of the window function; in pixels. 3
FacetClumps.KBins The coefficient used to calculate the number of eigenvalue bins. 35
FacetClumps.FwhmBeam The FWHM of the instrument beam; in pixels. 2
FacetClumps.VeloRes The velocity resolution of the instrument; in channels. 2
@@ Appendix B Catalogs of MWISP Clumps
The comprehensive catalogs of MWISP clumps are organized across three tables: detection parameters in Table 3, distance measurements in Table 4, and derived physical properties in Table 5. Examples presented include representative clumps with various distance determinations KDA_M as shown in Table 4, along with the selected clumps illustrated in Figure 16. The complete catalogs, containing all identified clumps and their parameters, are available through the online repository at DOI: 10.57760/sciencedb.28751.
Table 3: The Detection Catalog of Clumps.
@@
  • Note. Column (1): MWISP identification number; Columns (2)–(4): Position of the peak intensity in Galactic coordinates and velocity (Peak1, Peak2, Peak3); Columns (5)–(7): Centroid position in Galactic coordinates and velocity (Cen1, Cen2, Cen3); Columns (8)–(10): dispersion along the three axes (Size1, Size2, Size3); Column (11): Peak intensity in K; Column (12): Integrated intensity in K km s-1; Column (13): Volume of the clump in pixels; Column (14): Position angle of the major axis; Column (15): Edge flag (0: not at the edge of the survey, 1: at the edge); Column (16): ID of the SR to which the clump belongs.
Table 4: The Distance Catalog of Clumps.
MWISP CID GLON GLAT VLSR Dist Dist_EL Dist_EU KDA_S KDA_M KDA_Ref KDA_Flag
@@
  • Note. Column (1): MWISP identification number; Column (2): Galactic longitude in degrees; Column (3): Galactic latitude in degrees; Column (4): LSR velocity in km s-1; Column (5): Heliocentric distance in kpc; Columns (6)–(7): Lower and Upper error bounds on the distance estimate in kpc; Column (8): KDA solution; Column (9): Method used for distance determination; Column (10): References for distance determination. Column (11): KDA flag.
@@
    c. KDA_M: 0–clumps in SGQ and TGQ, 1-7–methods numbered as Figure 4.
    d. When KDA_M = 0 and 1, KDA_Ref is the references of maser: (1)(Reid et al., 2019), (2) (VERA Collaboration et al., 2020), (3) (Xu et al., 2021), (4) (Bian et al., 2022), (5) (Sakai et al., 2022), (6) (Li et al., 2022), and (7) (Bian et al., 2024). When KDA_M = 3, 5, and 7, KDA_Ref is the references of HI survey: (1) (HI4PI Collaboration et al., 2016), (2) (McClure-Griffiths et al., 2005), (3) (Taylor et al., 2003), (4) (Stil et al., 2006), and (5) (Peek et al., 2011). When KDA_M = 6, KDA_Ref is the references of IRDC: (1) (Simon et al., 2006), and (2) (Peretto & Fuller, 2009).
    e. When KDA_M = 0 and 1, KDA_Flag is the flag of maser-associated clumps: (1) maser-hosting, (2) maser-closing, and (3) neighbor of maser-closing.
Table 5: The Physical Parameter Catalog of Clumps.
MWISP CID GLON GLAT VLSR Dist Vfwhm Tex N(13CO) $R_{eq}$ Mass $\alpha_{vir}$ $\Sigma$ n(H2) $\Delta$ Flux $\Delta\sigma_{v}$ $\Delta R_{eq}$
(deg) (deg) (km s-1) (kpc) (km s-1) (K) (cm-2) (pc) (M) (M pc-2) (cm-3)
@@
  • Note. Column (1): MWISP identification number; Column (2): Galactic longitude in degrees; Column (3): Galactic latitude in degrees; Column (4): LSR velocity in km s-1; Column (5): Distance from the Sun in kpc; Column (6): FWHM velocity dispersion in km s-1; Column (7): Excitation temperature in K; Column (8): 13CO column density in cm-2; Column (9): Equivalent radius in pc; Column (10): Mass in solar masses; Column (11): Virial parameter; Column (12): H2 surface density in M pc-2. Column (13): H2 volume density in cm-3; Column (14): Flux fluctuation; Column (15): $\sigma_{v}$ fluctuation; Column (16): $R_{eq}$ fluctuation.

### v2

@@ 5.3.1 Cross-survey Comparison of Clump Scaling Relations
Figure 11 presents a detailed comparison of molecular clump scaling relations across catalogs, which can be directly assessed against Larson’s classical scaling laws (Larson, 1981), which have served as key references in molecular cloud dynamics.
The $M$ - $R_{\text{eq}}$ relation for molecular clumps (Figure 11a) exhibits a power-law index of $\alpha_{0.9}=2.229\pm 0.002$ ($P=0.97$), with CHIMPS ($\alpha_{0.9}=2.383$) and SEDIGISM ($\alpha_{0.9}=2.28$) displaying nearly identical slopes. These slopes, significantly steeper than the classical $M\propto R^{2.0}$ relation, indicate that surface density increases with size rather than remaining constant (Solomon et al., 1987; Lombardi et al., 2010). This departure from constant surface density spans multiple scales may reflect hierarchical structure formation, where material becomes increasingly concentrated at larger scales (Heitsch et al., 2009; Ballesteros-Paredes et al., 2011; Vázquez-Semadeni et al., 2017), potentially arising from the intrinsic structure of supersonic, intermittent turbulent flows (Roman-Duval et al., 2010).
@@
Figure 11(f) presents the $\sigma_{v}$ - $R_{\mathrm{eq}}\Sigma$ relationship, yielding a power-law index of $\alpha_{0.9}=0.303\pm 0.001$ with strong correlation ($P=0.83$) for MWISP clumps, paralleling earlier MWISP datasets (Feng et al., 2024). Both CHIMPS ($\alpha_{0.9}=0.383$, $P=0.69$) and SEDIGISM ($\alpha_{0.9}=0.412$, $P=0.76$) display slightly higher indices compared to MWISP. This relation demonstrates superior statistical robustness compared to the $\sigma_{v}$ - $R_{\mathrm{eq}}$ relation ($P=0.83$ versus $P=0.73$; (Heyer & Dame, 2015; Miville-Deschênes et al., 2017). For clumps in virial equilibrium, the virial theorem predicts $\sigma_{v}\propto(R\times\Sigma)^{0.5}$. The observed exponents remain below the virial prediction, indicating subvirial states where turbulent support scales more weakly with gravitational potential than expected in idealized equilibrium models.
While Larson’s original relations were based on relatively limited samples and the technologies available over 40 years ago (Larson, 1981), the comprehensive coverage and high precision of modern surveys like MWISP reveal nuances in molecular clump structure and dynamics that could not be captured previously (e.g. Solomon et al., 1987; Heyer et al., 2009; Lombardi et al., 2010; Ballesteros-Paredes et al., 2018). The MWISP survey’s unique characteristics, including large-scale coverage, an extensive sample, and uniform detection and distance analysis methods, make it well suited for analyzing correlations between parameters. These minimized systematic effects across diverse Galactic environments allow observed scaling relations to reflect more genuine physical processes. Rather than challenging the importance of Larson’s pioneering work, these results represent a natural advancement in our understanding of interstellar matter distribution enabled by modern observational capabilities.
Figure 12: Scaling relations between physical parameters of MWISP. Panel shows: (a) $M$-$R_{\rm eq}$, where light coral dotted lines indicate constant $\Sigma$ = 1, 10, 40, 100, 1000 Mpc-2 and lime dotted lines indicate constant $n({\rm H_{2}})$ = 10, 100, 1000, 10,000 cm-3; (b) $n({\rm H_{2}})$-$R_{\rm eq}$; (c) $\alpha_{\rm vir}$-$R_{\rm eq}$; (d) $\sigma_{v}$-$R_{\rm eq}$; (e) $\sigma_{v}$-$M$; (f) $\sigma_{v}$-$R_{\rm eq}\Sigma$; (g) $N(^{13}\mathrm{CO})$-$R_{\rm eq}$; (h) $\alpha_{\rm vir}$-$N(^{13}\mathrm{CO})$; and (i) $\sigma_{v}$-$N(^{13}\mathrm{CO})$. The distributions are shown for two subsets: MWISP complete clumps (gray points with cyan contours) and MWISP CMDs. Among the CMD subset, CMD-C, indicated by colored points with blue contours, represents clumps in the close vicinity of masers, while CMD-H, marked by red contours, represents clumps that host masers. For each relation, the best-fit power-law index and corresponding $P$ are indicated in the legend.
@@ 5.3.2 Impact of Star Formation Activity on Clump Scaling Relations
Figure 12 contrasts scaling relations between the complete MWISP sample and a subset of clumps with maser-determined distances (CMDs), further subdivided into clumps close to maser emissions (CMD-C, 1566 sources) and clumps directly hosting masers (CMD-H, 152 sources). These subsets represent different levels of star formation activity: CMD-H captures clumps with ongoing maser excitation processes indicative of active star formation (e.g., Menten, 1991; Cragg et al., 2005; Breen et al., 2010; Urquhart et al., 2011b), CMD-C represents regions influenced by nearby star formation activity, while the complete sample encompasses diverse star formation environments ranging from quiescent to active phases.
@@
The $\alpha_{\rm vir}$-$R_{\rm eq}$ relation (Figure 12c) displays negative correlations, with changes in slope from the complete sample ($\alpha_{0.9}=-0.728$) to CMD-C ($\alpha_{0.9}=-0.69$, P=0.56P=-0.56) to CMD-H ($\alpha_{0.9}=-0.582$, P=0.59P=-0.59). This progressive flattening of the power-law slope, combined with the rightward and downward displacement of CMD-H distribution, indicates that maser-hosting clumps are systematically larger and more gravitationally bound. The flattening trend suggests that virial parameters become less sensitive to radius variations in active star formation environments, possibly reflecting stellar feedback processes that provide scale-independent support mechanisms across different clump sizes, partially balancing gravitational binding with nonthermal support to maintain relatively uniform dynamical states in these regions.
Panels (d)-(f) in Figure 12 demonstrate changes in velocity dispersion scaling relations across different star formation environments. In the $\sigma_{v}$$R_{\rm eq}$ relation (panel (d)), both maser-associated populations display flatter slopes ($\alpha_{0.9}=0.152,0.161$, $P=0.27,0.41$) compared to the complete sample ($\alpha_{0.9}=0.396$). This pattern continues in the $\sigma_{v}$-$M$ relation (panel (e)), where power-law indices decrease from the complete sample ($\alpha_{0.9}=0.172$) to CMD-C ($\alpha_{0.9}=0.151$, $P=0.63$) and CMD-H ($\alpha_{0.9}=0.124$, $P=0.55$). Similarly, the $\sigma_{v}$-$R_{\rm eq}\Sigma$ relation (panel (f)) shows a decline from the complete sample ($\alpha_{0.9}=0.303$) to CMD-C ($\alpha_{0.9}=0.287$, $P=0.72$) and CMD-H ($\alpha_{0.9}=0.261$, $P=0.74$). The overall flattening of these scaling relations indicates that velocity dispersion becomes less sensitive to physical parameters in active star formation environments. The relatively small differences between CMD-C and CMD-H coefficients suggest that these modifications to turbulent properties occur in association with star formation activity, irrespective of whether clumps directly host masers.
@@
The $\sigma_{v}$-$N(^{13}\mathrm{CO})$ relation (Figure 12i) shows substantial reduction in slopes from the complete sample ($\alpha_{0.9}=0.801$, $P=0.82$) to identical values for both CMD-C ($\alpha_{0.9}=0.434$, $P=0.63$) and CMD-H ($\alpha_{0.9}=0.429$, $P=0.72$). This relation provides particularly valuable insights as both parameters are distance-measurement-independent quantities, eliminating potential biases from distance uncertainties. The power-law indices demonstrate higher-density regions in general clumps maintain proportionally higher turbulent motions. The identical slopes between CMD-C and CMD-H populations indicate that these relationships converge to similar characteristics in maser-associated environments, suggesting that turbulent properties may be regulated by similar physical processes (Heyer & Brunt, 2004) in maser-associated regions.
@@ 6 Summary
We present an equivalent global detection scheme that significantly reduces computational resource requirements while preserving detection integrity for processing large astronomical datasets. By integrating this scheme with FacetClumps, we successfully extract 71,661 molecular clumps across the MWISP Phase I survey’s high-resolution 13CO data spanning 2310 deg2. The designed automatic hierarchical distance decision method, using SRs as fundamental objects, effectively resolves the KDA problem, offering reliable distance measurements for 97.94% of clumps.
The MWISP clumps have median values of 225.78 $M_{\odot}$ in mass, 1.72 pc in radius, 0.4 km s-1 in velocity dispersion, 25.53 $M_{\odot}$ pc-2 in surface density, and 214.69 cm-3 in volume density. With a median virial parameter of $\alpha_{\text{vir}}=1.44$, these structures are predominantly gravitationally bound, as 65.3% satisfy the criterion $\alpha_{\text{vir}}<2$ and account for 96.3% of the total statistical mass. The MWISP clump catalog uniquely traces intermediate-mass and intermediate-size structures with narrow velocity dispersions and low densities. These clumps, characterized by extended, diffuse molecular distributions, represent an important component of the Galactic molecular clump population.
@@
    The 13CO column density - radius relation ($\bar{N}(^{13}\mathrm{CO})\propto R^{0.296}$) shows a gradual column density increase with clump size.
@@
The MWISP clump catalogs feature broad coverage, effectively capturing the critical transition from diffuse molecular gas to gravitationally collapsing structures. Despite intrinsic survey limitations, completeness limits, and distance effects, these catalogs achieve exceptional statistical robustness and represent the most extensive study of molecular clumps to date. This unique characteristic establishes these catalogs the basis for a series of investigations, including: (1) analyzing the distributions of clump parameters across the Galactic disk and their trends with Galactocentric distance (Sun et al., 2024b); (2) identifying filaments using clump masks, directional properties, and positional properties (Jiang et al., 2025); (3) integrating multiwavelength observations to accurately classify the evolutionary stages of star-forming regions (Urquhart et al., 2022) and derive their associated scaling relations; and numerous additional applications. These investigations will significantly advance our understanding of the complex interplay between molecular clumps, star formation, and Galactic evolution.
@@ Acknowledgements
We are grateful to the anonymous referee for the invaluable insights and comments, which enabled us to refine and enhance this work. This work is supported by the National Key R&D Program of China (grant No. 2023YFA1608001) and the National Natural Science Foundation of China (grant No. U2031202). This research makes use of the data from the Milky Way Imaging Scroll Painting (MWISP) project, which is a multiline survey in 12CO/13CO/C18O along the northern Galactic plane with the PMO-13.7m telescope. We are grateful to all the members of the MWISP working group, particularly the staff members at PMO 13.7m telescope, for their long-term support. MWISP was sponsored by the National Key R&D Program of China with grants 2023YFA1608000, 2017YFA0402701, and the CAS Key Research Program of Frontier Sciences with grant QYZDJSSW-SLH047.
@@ Appendix A Configuration Parameters
Table A.1: Parameters of FacetClumps along with their explanations and values.
@@
FacetClumps.RMS The noise rms or S/N of the data 1 (S/N data)
@@
FacetClumps.SWindow The scale of the window function; in pixels 3
FacetClumps.KBins The coefficient used to calculate the number of eigenvalue bins 35
FacetClumps.FwhmBeam The FWHM of the instrument beam; in pixels 2
FacetClumps.VeloRes The velocity resolution of the instrument; in channels 2
@@ Appendix B Catalogs of MWISP Clumps
The comprehensive catalogs of MWISP clumps are organized across three tables: detection parameters in Table B.1, distance measurements in Table B.2, and derived physical properties in Table B.3. Examples presented include representative clumps with various distance determinations KDA_M as shown in Table B.2, along with the selected clumps illustrated in Figure E.1. The complete catalogs, containing all identified clumps and their parameters, are available through the online repository at DOI: 10.57760/sciencedb.28751.
Table B.1: The Detection Catalog of Clumps.
@@
  • Note. Column (1): MWISP identification number. Columns (2)–(4): position of the peak intensity in Galactic coordinates and velocity (Peak1, Peak2, and Peak3). Columns (5)–(7): centroid position in Galactic coordinates and velocity (Cen1, Cen2, and Cen3). Columns (8)–(10): dispersion along the three axes (Size1, Size2, and Size3). Column (11): peak intensity in K. Column (12): Integrated intensity in K km s-1. Column (13): volume of the clump in pixels. Column (14): position angle of the major axis. Column (15): edge flag (0: not at the edge of the survey; 1: at the edge). Column (16): ID of the SR to which the clump belongs.
Table B.2: The Distance Catalog of Clumps.
MWISP CID GLon GLat $V_{LSR}$ Dist Dist_EL Dist_EU KDA_S KDA_M KDA_Ref KDA_Flag
@@
  • Note. Column (1): MWISP identification number. Column (2): Galactic longitude in degrees. Column (3): Galactic latitude in degrees. Column (4): LSR velocity in km s-1. Column (5): heliocentric distance in kpc. Columns (6)–(7): lower and upper error bounds on the distance estimate in kpc. Column (8): KDA solution. Column (9): method used for distance determination. Column (10): references for distance determination. Column (11): KDA flag.
@@
    c. KDA_M: 0–clumps in SGQ or TGQ, 1-7–methods numbered as Figure 4.
    d. When KDA_M = 0 or 1, KDA_Ref is the references of the maser: (1)(Reid et al., 2019), (2) (VERA Collaboration et al., 2020), (3) (Xu et al., 2021), (4) (Bian et al., 2022), (5) (Sakai et al., 2022), (6) (Li et al., 2022), and (7) (Bian et al., 2024). When KDA_M = 3, 5, or 7, KDA_Ref is the references of the HI survey: (1) (HI4PI Collaboration et al., 2016), (2) (McClure-Griffiths et al., 2005), (3) (Taylor et al., 2003), (4) (Stil et al., 2006), and (5) (Peek et al., 2011). When KDA_M = 6, KDA_Ref is the references of the IRDC: (1) (Simon et al., 2006), and (2) (Peretto & Fuller, 2009).
    e. When KDA_M = 0 or 1, KDA_Flag is the flag of maser-associated clumps: (1) maser-hosting, (2) maser-closing, and (3) neighbor of maser-closing clump.
Table B.3: The Physical Parameter Catalog of Clumps.
MWISP CID GLon GLat $V_{LSR}$ Dist $V_{fwhm}$ $T_{ex}$ N(13CO) $R_{eq}$ Mass $\alpha_{vir}$ $\Sigma$ nn(H2) $\Delta$ Flux $\Delta\sigma_{v}$ $\Delta R_{eq}$
(deg) (deg) (km s-1) (kpc) (km s-1) (K) (cm-2) (pc) ($M_{\odot}$) ($M_{\odot}$ pc-2) (cm-3)
@@
  • Note. Column (1): MWISP identification number. Column (2): Galactic longitude in degrees. Column (3): Galactic latitude in degrees. Column (4): LSR velocity in km s-1. Column (5): distance from the Sun in kpc. Column (6): FWHM velocity dispersion in km s-1. Column (7): excitation temperature in kelvins. Column (8): 13CO column density in cm-2. Column (9): equivalent radius in pc. Column (10): mass in solar masses; Column (11): virial parameter. Column (12): H2 surface density in M pc-2. Column (13): H2 volume density in cm-3. Column (14): flux fluctuation; Column (15): $\sigma_{v}$ fluctuation. Column (16): $R_{eq}$ fluctuation.